OpenStudy (anonymous):

if f (x)=2x-x^2+1, find (x-3)

OpenStudy (anonymous):

\[f(x)=2x-x^2+1\] \[f(x-3)=2(x-3)-(x-3)^2+1\] then algebra

OpenStudy (anonymous):

I dont understand

OpenStudy (anonymous):

what you need to do is plug x-3 in x's in the f(x)! and solve

OpenStudy (anonymous):

f(x)=2x-x^2+1 plug x-3 in f(x) f(x-3)=2(x-3)-(x-3)^2+1 f(x-3)=2x-6-x^2+6x-9+1 f(x-3)=8x-14-x^2

OpenStudy (anonymous):

Thank you! Got it! :)

OpenStudy (anonymous):

u r welcome..